OpenStudy (anonymous):

if x=.28 determine the three real solutions (zeros) of F(X)=x^5-7x^4+10x^3+14x^2-21x

OpenStudy (anonymous):

What does it mean by if x=0.28?

OpenStudy (anonymous):

it's a function and .28 has to be plugged in where there is an X

OpenStudy (anonymous):

Okay x^5-7x^4+10x^3+14x^2-21x=0 x(x^4-7x^3+10x^2+14x-21)=0 So one solution is x=0. If we divide through by x we are left with x^4-7x^3+10x^2+14x-21=0 which has two complex solutions and two real ones.

OpenStudy (anonymous):

i did not get the last part???

OpenStudy (anonymous):

I just took the x to the other side so I would be left with x^4-7x^3+10x^2+14x-21=0 I check wolfram and the are two real solutions and two complex one. http://www.wolframalpha.com/input/?i=x^5-7x^4%2B10x^3%2B14x^2-21x These other solutions we can estimate but not find exact answers.

OpenStudy (anonymous):

What grade/ level are you in?

OpenStudy (anonymous):

i am in high school senior year

OpenStudy (anonymous):

Okay good, one way we can estimate is by graphing it on your graphics calculator and have it estimate the other two solutions. Another way is to write it up in a program and use Newton's Method or another root finding tool.

OpenStudy (anonymous):

yeah i have a graphing calc so how would i do it? if u can help me!!! please

OpenStudy (anonymous):

Okay first off we need to graph y=x^5-7x^4+10x^3+14x^2-21x, are you able to do this? What model/brand is your graphics calculator?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

TI-84 texas

OpenStudy (anonymous):

okay i graphed it

OpenStudy (anonymous):

Good, that's the calculator I had in high school too. Makes this easier.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do i do now?

OpenStudy (anonymous):

Now press 2nd Trace

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't have the calculator anymore, but there should be some option called zeros or roots. Choose that option

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it says lef bound? right bound?

OpenStudy (anonymous):

good, now for the left bound you want to choose the closest point you can get before the graph crosses the x axis for the right bound you want to get it as close to x axis after you've crossed it.

OpenStudy (anonymous):

Remeber you have to do this step for each point the graph crosses the x-axis. So for your function you need to do this 3 times to get all the real roots.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What answers are you getting?

OpenStudy (anonymous):

-1.36823 and 1.18796

OpenStudy (anonymous):

Good work, that's correct. :D

OpenStudy (anonymous):

yay thanks to u